(1 point) Solve the equation 2х — 2у + 7z %3D0 х y | = s
Answer #1

2x - R9 117 - 0 31 - 34: - 7 x-y =772 108 x = 0, y =77 $ 2-2 fort X = -7 y=0 $.2=2 (0,2,2) and (-7,0,2) 20€ two linearly inde